The tallest living man at one time had a height of 231 cm. The shortest living man at that time had a height of 142.1 cm. Heights of men at that time had a mean of 172.16 cm and a standard deviation of 5.22 cm. Which of these two men had the height that was more​ extreme?


1
Expert's answer
2022-05-18T06:51:43-0400
"z_1=\dfrac{231-172.16}{5.22}=11.2720"

"z_2=\dfrac{142.1-172.16}{5.22}=-5.7586"

Since the z score for the tallest man is "z_1=11.2720" and the z score for the shortest man is "z_2=-5.7586," "|z_1|=11.2720>5.7586=|z_2|," the tallest man had the height that was more extreme.